Hello, I would like to get help with the following algorithms and their respective analysis of runtime along with their recurrence equations, thanks in advance.

1.       Analyze each of the following algorithms by providing a tight big-Oh bound on the run time with respect to the value n. Create a recurrence equation.

a.    

void padawan(int n)

{

      if( n <= 10 )

            time++;

      else

      {

            for(int i=0; i<n; i++)

                        time++;

            padawan(n/3);

            padawan(n/3);

            padawan(n/3);

      }

}

b.   

void nerfHerder(int n)

{

      if( n <= 15 )

            time++;

      else

      {

            for(int i=0; i<n; i++)

                  time++;

            nerfHerder(n-1);

      }

}

c.   

void squire(int n)

{

      if( n <= 12 )

            time++;

      else

      {

            squire(n/2);

            squire(n/2);

            for(int i=0; i<n; i++)

                  for(int j=i; j<n; j++)

                        time++;

            squire(n/2);

            squire(n/2);

            squire(n/2);

            squire(n/2);

            squire(n/2);

            squire(n/2);

      }

}
Answer #1
Master method

T(n) = aT(n/b) + f(n) where a>=1 and b>1

There are following three cases:
1. If f(n) = Θ(nc) where c < Logba then T(n) = Θ(nLogba)

2. If f(n) = Θ(nc) where c = Logba then T(n) = Θ(ncLog n)

3.If f(n) = Θ(nc) where c > Logba then T(n) = Θ(f(n))

a) Given equation runtime

Runtime for if condition is O(1)

In else condition runtime for loop is O(n) and for recurrence relation runtime can be written as

 

By using Master method as as c=logab it falls under case 2, T(n)=

b)runtime for loop is O(n)

recurrence relation is

and c>logab as c=1,a=1,b=1 it falls under case 3,T(n)=

3)runtime for for loops is O(n2)

As there are 6 recursive functions

c<logab because a=6,b=2 and c=2 this falls under case 1,Therefore